OpenStudy (anonymous):

What is the vertex of the parabola given by y = x2 − 10x + 17?PLEASE HELP ME!!!!!!!!!

OpenStudy (anonymous):

please help me :))

OpenStudy (anonymous):

its 2 points. x and y

OpenStudy (anonymous):

welll....

OpenStudy (anonymous):

have you solvedd the problem yet???

OpenStudy (amistre64):

yeah lucas; havent you solved it yet lol

OpenStudy (anonymous):

cant YOU solve it amistre64?

OpenStudy (anonymous):

The vertex is at (5,8)

OpenStudy (anonymous):

omg ppl!

OpenStudy (anonymous):

Parabola is in standard form. Put into vertex form. y(x)+25=x^2-10x+25+17 y(x)+25=x^2-10x+25+17 y(x)+25= (x-5)^2 + 17 y(x)= (x-5)^2 -8 Vertex form formula is y= a(x-h)2+k -h,k is coordinate when in vertex form. (5,-8)

OpenStudy (amistre64):

the question is: cant YOU solve it jnash? just cause the site is free doesnt mean we have to do YOUR work FOR you lol

OpenStudy (anonymous):

yes, my bad (5,-8)

OpenStudy (anonymous):

lol ;) i know lol i just need help you guys ;P

OpenStudy (anonymous):

needed*

OpenStudy (amistre64):

:) youll do fine ;)

OpenStudy (anonymous):

Which one is a complex number in standard form for the expression (1 − 3i) − (5 + 4i) − 12i?

OpenStudy (anonymous):

;)